Deadeye:

Solve this problem on paper using all four steps. A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week? They sold boxes the first week, boxes the second week, and boxes the third week.

Hoodmemes:

OK so x is first week,2nd week is y,3rd is z

Hoodmemes:

5 more the 2nd week than 1st week = x+5=y

Hoodmemes:

And 3rd week is double 2nd week=z=2y

Hoodmemes:

trx+y+z=431 sub,y=5+x and z=2y,x+5+x+2y=431,sub 5+x for y,x+5+x+2(5+x)=431,x+5+x+10+2x=431

Hoodmemes:

Then it would be minus 15, 4x=41,divide by 4 , x=104, add 4,x+5=109

Hoodmemes:

Ight so now double 2*109=218

Hoodmemes:

SO now u would get your answer, and start it off wit they sold,1st week: (blank),2nd week: (blank),and 3rd week: (blank)